# Semiparametric Discovery and Estimation of Interaction in Mixed Exposures using Stochastic Interventions

### v2

@@ 6.4 Methods Applied to NHANES Dataset
The National Health and Nutrition Examination Survey (NHANES) 2001-2002 cycle serves as the foundation of our analysis. This data source, known for its credibility in the public health domain, boasts interviews with 11,039 individuals. Of this subset, 4,260 provided blood samples and willingly consented to DNA analysis. The data set used aligns with that of Mitro et al. Mitro et al. [2016] focusing on the correlation between persistent exposure to organic pollutants (POPs), specifically those binding to the aryl hydrocarbon receptor (AhR) and extended leukocyte telomere length (LTL). However, this subset was further refined to ensure complete exposure data, yielding 1007 participants for our analysis, compared to 1003 in the initial study by Mitro et al. In alignment with protocols detailed by Mitro et al. Mitro et al. [2016], exposure was quantified, focusing on 18 congeners. These include 8 non-dioxin-like PBCs, 2 non-ortho PCBs, 1 mono-ortho PCB, 4 Dioxins, and 4 Furans. All congeners underwent lipid serum adjustments using an enzymatic summation methodology. The telomere length, was analyzed using the quantitative polymerase chain reaction (qPCR) methodology Mitro et al. [2016]. This technique measures the T/S ratio by comparing the length of the telomere to a standardized reference DNA. To enhance accuracy, samples underwent triple assays in duplicate, generating an average from six data points. The CDC conducted a blinded quality control assessment. Our modeling accounted for several covariates, including demographic factors such as age, sex, race / ethnicity and level of education, as well as health indicators such as BMI, serum cotinine, and blood cell distribution and count. Categorizations for race/ethnicity, education, and BMI are consistent with previous studies. Furthermore, this comprehensive dataset is integrated into the SuperNOVA package for replicable analysis.
@@
We configured our methodology using default learners and employed a 20-fold CV without a quantile threshold for the F-statistic of the basis function. The built-in algorithms of the Super Learner were used to identify variable associations in the discovery procedure, using linear terms and two-way interactions of earth models and various hyperparameters (13 models in total). Given the data range, a $\delta$ value of 2 was fixed for all exposures, denoting a focus on counterfactual changes in telomere length for an increase in 2 ng / g or pg / g across exposures that predict telomere length depending on the scale of exposure.
6.5 Comparing NHANES Furan Results to Previous Work
In our analysis of the 18 POPs, furan 2,3,4,7,8-pncdf emerged consistently across all 20 folds. None of the other chemicals were featured in any fold. The pooled TMLE estimated a 0.02 (0.03 - 0.001) decrease in telomere length after a 2 pg / g increase in furan 2,3,4,7,8-pncd which was significant at p = 0.04. Results in individual folds showed similar effects. The Mitro et al. [2016] summarize the results of the workshop. They state that clustering methods identified high, medium, and low POP exposure groups with longer log-LTL observed in the high exposure group. Principal component analysis (PCA) and exploratory factor analysis (EFA) revealed positive associations between overall POP exposure and specific POPs with log-LTL. Penalized regression methods identified three congeners (PCB 126, PCB 118, and furan 2,3,4,7,8-pncdf) as potentially toxic agents. WQS identified six POPs (furans 1,2,3,4,6,7,8-hxcdf, 2,3,4,7,8-pncdf, and 1,2,3,6,7,8-hxcdf, and PCBs 99, 126, 169) as potentially toxic agents with a positive overall effect of the POP mixture. BKMR found a positive linear association with furan 2,3,4,7,8-pncdf, suggestive evidence of linear associations with PCBs 126 and 169, a positive overall effect of the mixture, but no interactions among congeners. These results (in the supervised methods) controlled for the same covariates. Interestingly, although we corroborate the finding of furan 2,3,4,7,8-pncdf, specifically in the BKMR method (the most flexible method used), our results show a negative association. Additionally, we do not find associations of any other POP used in any fold, meaning that in the best-fitting spline model, no basis functions for any of the other POPs were used. The inverse association we find compared to these other methods could be a result of our $Q$ model being a Super Learner, which can model relationships perhaps missed in these other methods. Through this NHANES example, we underscore the ability of SuperNOVA in discerning pertinent variable relationships in complex mixture data, subsequently producing estimates. The consolidated findings allow for a clear understanding of the potential changes in outcomes in different exposure scenarios, improving the interpretability of the results. Interestingly, despite the heterogeneity of the findings of other methods, we identify only one chemical possibly related to a reduced telomere length, furan 2,3,4,7,8-pncdf, which warrants further investigation.
@@ 7 Discussion
In our study, we developed a semiparametric statistical model to estimate interaction effects in mixed exposures, focusing on identifying and quantifying exposure sets that influence outcomes. Using data-adaptive target parameters and using targeted maximum likelihood estimation (TMLE) within a cross-validation framework, we achieved asymptotically unbiased estimates of interaction effects. Our approach offers several advantages, including robustness to the number of exposures and the complexity of the data-generating process, and it provides interpretable results that improve understanding of drug synergies and environmental health impacts. Unlike traditional interaction analysis methods, our model minimizes assumptions, yielding more reliable and interpretable estimates of interaction effects and focuses on causal quantities. Our methodology faces two primary limitations. First, the computational intensity of the density estimation step, though mitigated by parallel processing, remains a challenge. Alternatives such as recasting density estimation as a classification problem could offer efficiency improvements, but would require iterative refitting of the classifier. Second, the identification of exposure sets varies between cross-validation folds, introducing variability that demands careful and consistent reporting to ensure reproducibility and reliability of results. Future enhancements, including the incorporation of algorithms like glmnet and the highly adaptive lasso (HAL), are expected to stabilize identification processes. It is crucial for findings to be reported only if they demonstrate consistency across multiple cross-validation folds, with transparent communication about the variability of these findings.
In future research, we aim to address the limitations and expand the capabilities of our approach. Key directions include improving density estimation. We plan to explore the transformation of density estimation into a classification problem using Bayesian probability. This shift promises to improve computational efficiency and incorporate a wider range of machine learning algorithms. Refining pooling strategies and testing is also required; we will develop and evaluate new pooling strategies, such as kernel estimators to estimate the sampling distribution of our estimates with null effects and sampling from null distributions to refine pooled estimates incorporating the null, hopefully improving the robustness of our inverse variance strategy. We can also focus beyond two-way interactions to include higher-order complexities. We also intend to investigate alternative TMLE strategies that may offer better performance, especially in smaller samples. This includes direct computation of influence functions for interaction parameters and novel estimation techniques for clever covariates. To facilitate adoption, we have developed the SuperNOVA R package, available on GitHub, offering researchers software to apply our proposed methodology for the analysis of mixtures, this package has been peer reviewed by the Journal of Open-Source Software McCoy et al. [2023].

### v3

@@ 7.3 Methods Applied to NHANES Dataset
The National Health and Nutrition Examination Survey (NHANES) 2001-2002 cycle serves as the foundation of our analysis. This data source, known for its credibility in the public health domain, boasts interviews with 11,039 individuals. Of this subset, 4,260 provided blood samples and willingly consented to DNA analysis. The data set used aligns with that of Mitro et al. Mitro et al. [2016] focusing on the correlation between persistent exposure to organic pollutants (POPs), specifically those binding to the aryl hydrocarbon receptor (AhR) and extended leukocyte telomere length (LTL). However, this subset was further refined to ensure complete exposure data, yielding 1007 participants for our analysis, compared to 1003 in the initial study by Mitro et al. In alignment with protocols detailed by Mitro et al. Mitro et al. [2016], exposure was quantified, focusing on 18 congeners. These include 8 non-dioxin-like PBCs, 2 non-ortho PCBs, 1 mono-ortho PCB, 4 Dioxins, and 4 Furans. All congeners underwent lipid serum adjustments using an enzymatic summation methodology. The telomere length, was analyzed using the quantitative polymerase chain reaction (qPCR) methodology Mitro et al. [2016]. This technique measures the T/S ratio by comparing the length of the telomere to a standardized reference DNA. To enhance accuracy, samples underwent triple assays in duplicate, generating an average from six data points. The CDC conducted a blinded quality control assessment. Our modeling accounted for several covariates, including demographic factors such as age, sex, race / ethnicity and level of education, as well as health indicators such as BMI, serum cotinine, and blood cell distribution and count. Categorizations for race/ethnicity, education, and BMI are consistent with previous studies. Furthermore, this comprehensive dataset is integrated into the InterXshift package for replicable analysis.
Gibson et al. [2019] reanalyzed this same data using more contemporary statistical methods. They found that clustering methods identified high, medium, and low POP exposure groups with longer log-LTL observed in the high exposure group. Principal component analysis (PCA) and exploratory factor analysis (EFA) revealed positive associations between overall POP exposure and specific POPs with log-LTL. Penalized regression methods identified three congeners (PCB 126, PCB 118, and furan 2,3,4,7,8-pncdf) as potentially toxic agents. WQS identified six POPs (furans 1,2,3,4,6,7,8-hxcdf, 2,3,4,7,8-pncdf, and 1,2,3,6,7,8-hxcdf, and PCBs 99, 126, 169) as potentially toxic agents with a positive overall effect of the POP mixture. BKMR found a positive linear association with furan 2,3,4,7,8-pncdf, suggestive evidence of linear associations with PCBs 126 and 169, a positive overall effect of the mixture, but no interactions among congeners. These results (in the supervised methods) controlled for the same covariates. Figure 2 shows the results for WQS, BKMR and Lasso regression from their paper. We highlight furan 2,3,4,7,8-pncdf as the chemical which shows consistent positive association with LTL across the methods. As such, we might expect that this chemical would consistently be found as the top positive marginal impact based on our target parameter across all the folds.
Figure 2: Mixture Methods Results from Gibson et. al.
@@
We configured our methodology using default learners and employed a 5-fold CV. Given the data range, a $\delta$ value of half a standard deviation was used for all exposures, denoting a focus on counterfactual changes in telomere length for an increase in half a standard deviation across exposures that predict the telomere length depending on the scale of exposure.
7.4 NHANES Furan Results
Generally, when applying our InterXshift method to identify the top positive and negative marginal effects; and the top synergistic and antagonistic interactions, we do not find consistent estimates for these parameters. Table 8 shows the estimates of the top positive association between the folds. Although we see 2,3,4,7,8-pncdf is identified as the top positive effect in 3 of the folds, 2 of the folds estimate a different chemical as the top positive effect; therefore, the estimates are inconsistent. Likewise, the estimates for the top inverse association with LTL were also inconsistent across the folds. The top synergistic and antagonistic interactions were also inconsistent and, therefore, indicate no interaction in the mixed POP exposure.
Condition Psi Variance SE Lower CI Upper CI P-value Fold N Delta
1 2,3,4,7,8-pncdf 0.02 0.00 0.02 -0.02 0.05 0.42 1 202 2.88
2 2,3,4,7,8-pncdf 0.01 0.00 0.02 -0.03 0.04 0.64 2 202 2.88
3 PCB74 -0.05 0.00 0.03 -0.11 0.02 0.18 3 201 6794.89
4 1,2,3,4,7,8-hxcdf 0.00 0.00 0.02 -0.04 0.05 0.92 4 201 2.49
5 2,3,4,7,8-pncdf 0.03 0.00 0.02 -0.01 0.08 0.12 5 201 2.88
6 Rank 1 -0.00 0.00 0.01 -0.03 0.02 0.74 Pooled TMLE 1007 1361.20
Table 8: Rank 1 Results for Different Conditions
@@ 8 Discussion
In our study, we developed a semiparametric statistical model to discover and estimate interaction effects in mixed exposures. Using data-adaptive target parameters paired with TMLE within a cross-validation framework, we achieved asymptotically unbiased estimates of identifying and estimating the top interaction effects in a mixed exposure. Our approach offers several advantages, including robustness to the number of exposures and the complexity of the data-generating process, and provides interpretable results that improve understanding of drug synergies and environmental health impacts. Unlike traditional interaction analysis methods, our model minimizes assumptions, yielding more reliable and interpretable estimates of interaction effects, and focuses on causal quantities. Our methodology faces the primary limitation that the identification of exposure sets that are estimated for the top marginal, synergistic, and antagonistic interaction may vary between cross-validation folds, introducing variability that demands careful and consistent reporting to ensure reproducibility and reliability of results. However, rather than a complete limitation, this indicates the strength of the association in subsets of the data. If the “signal“ is consistent, then the findings for each parameter will be consistent.
In future research, our aim is to improve the estimation of density estimation ratios and to evaluate the variability of estimates using various methods for estimation of density ratios. We can also focus beyond two-way interactions to include higher-order complexities. We also intend to investigate alternative TMLE strategies that may offer better performance, especially in smaller samples. To facilitate adoption, we have developed the InterXshift R package, available on GitHub, offering researchers software to apply our proposed methodology for the analysis of mixtures, this package has been peer reviewed by the Journal of Open-Source Software McCoy et al. [2023].